OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. If p is the number of pounds of jelly beans Terri purchases, what expression could be used to represent the number of pounds of chocolate drops she purchases? (Points : 1) 2.1 – p 2.1p p + 2.1 p – 2.1

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

i believe it's D

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

more not alot

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. If p = pounds of caramels, which equation in simplest form represents the situation described. (Points : 1) –0.16p + 1.62 = 1.48 –0.16p + 1.62= 1.7 0.16p + 1.34 = 1.48 0.16p + 1.34 = 1.7

OpenStudy (jonnyvonny):

Seems like you just want answers. I shall not partake.

OpenStudy (anonymous):

All - jellybeans = chocolate so 2.1 - p = x so it's A :) I'm 100% sure

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@JonnyVonny thats because I'm very behind and i'm really trying to get back up because iv been very very sick and alot has happened and im sorry

OpenStudy (jonnyvonny):

No need to apologize to me, you didn't harm me. Just make sure you learn how to do this, for, in higher math's you'll need to know how to do this already, prior to actually doing the problem.

OpenStudy (anonymous):

yeah your right @JonnyVonny can you help me

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. If p is the number of pounds of jelly beans Terri purchases, what expression could be used to represent the number of pounds of chocolate drops she purchases? (Points : 1) 2.1 – p 2.1p p + 2.1 p – 2.1

OpenStudy (anonymous):

so its either A or D

OpenStudy (anonymous):

i think its D again

OpenStudy (anonymous):

but I dont know which one to choose from

OpenStudy (mathmath333):

its A

OpenStudy (anonymous):

wait no its A

OpenStudy (mathmath333):

2.1 – p

OpenStudy (anonymous):

thank you A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. If p = pounds of caramels, which equation in simplest form represents the situation described. (Points : 1) –0.16p + 1.62 = 1.48 –0.16p + 1.62= 1.7 0.16p + 1.34 = 1.48 0.16p + 1.34 = 1.7

OpenStudy (mathmath333):

i think its C 0.16p + 1.34 = 1.48

OpenStudy (anonymous):

thank you A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. How many pounds of chocolate-covered pretzels and gumdrops did Troy buy? (Points : 1) 0.5 pounds of pretzels; 1.3 pounds of gumdrops 0.9 pounds of pretzels; 0.9 pounds of gumdrops 1.3 pounds of pretzels; 0.5 pounds of gumdrops 0.8 pounds of pretzels; 1 pound of gumdrops

OpenStudy (anonymous):

i have two more after this then I'm done :)

OpenStudy (mathmath333):

use this equations to get that \(\large\tt \color{black}{p+c=1.7}\) \(\large\tt \color{black}{0.95p+0.79=1.48}\)

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

again same procedure use this equations solve for p and g \(\large\tt \color{black}{p+g=1.8}\) \(\large\tt \color{black}{0.89p+0.65g=1.29}\)